OpenStudy (anonymous):

Fan! please help!! 2(3y + 7 + 5) = 196 − 16?

OpenStudy (kidrah69):

\(\huge\color{orange}{WELCOME~TO~OpenStudy!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (kidrah69):

Fist distribute :)

OpenStudy (kidrah69):

\[2(3y + 7 + 5) = 196 − 16? \] \[2(3y + 7 + 5)\]

OpenStudy (anonymous):

so now i have 5y + 14 + 10 = 196 - 16?

OpenStudy (kidrah69):

No u need to MULTIPLY not add remember we are distributing :)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i knew that, hold on lol

OpenStudy (anonymous):

so its 6y + 14 + 10 =196 - 16

OpenStudy (kidrah69):

yes! :) what is 14+10=? 196-16=?

OpenStudy (anonymous):

14 + 10 is 24 and 196 - 16 is 180

OpenStudy (kidrah69):

6y+40=180 and we are trying to find the value of Y right?

OpenStudy (anonymous):

yes, so i would subtract by 24 on both sides right?

OpenStudy (kidrah69):

lol you noticed my typo! wow thats amazing :0 yes you are correct :) 180-24=?

OpenStudy (anonymous):

so Y = 26!!

OpenStudy (anonymous):

2 (3 y + 7 + 5) = 196 - 16 24 + 6 y = 196 - 16 6 y = 196 - 16 - 24 6 y = 156 y = 26

OpenStudy (anonymous):

thanks!

OpenStudy (kidrah69):

=) that was right

OpenStudy (anonymous):

thank you!